OpenStudy (anonymous):

at the dog park there are several dogs with their owners. countin heads there are 20 , counting legs 62. how many dogs and owners are there?

OpenStudy (anonymous):

Let x be the number of owners and let y be the number of dogs. So for your first equation x + y = 20 For your second equation 2x + 4y = 62 Write x in terms of y x = 20 - y Substitute in the second equation 2(20 - y) + 4y = 62 ===> 40 - 2y + 4y = 62 2y + 40 = 62 2y = 22 y = 11 So there are 11 dogs and 9 owners